Use the formula to find the number of moles in


2.5 X 10 -6 dm3 of a 8.0 X 10 -6 moldm-3 solution of potassium   permanganate.



(a)           Evaluate showing your working out without calculator use:-     


           2.6x107 x 3x10-4      

                        1.3 x 103               


1
Expert's answer
2021-03-09T01:50:43-0500

(a) "2.5*10^{-6}*8*10^{-6}=20*10^{-12}=2*10^{-11}\; mol."


(b) "2.6*10^7*3*10^{-4}=7.8*10^3=7800."

"1.3*10^3=1300."